# Image-Based Virtual Try-On: A Survey

### v2

@@ 2.3 Cloth-agnostic Person Representation
RGB image $\mathcal{P}_{\begin{subarray}{c}1,2,3,4\end{subarray}}$ provides pixel-level features that can better preserve the identity information of individuals and the content outside the try-on area. $\mathcal{P}_{\begin{subarray}{c}1\end{subarray}}$ shows the original person image which is disturbed by the original clothes. $\mathcal{P}_{\begin{subarray}{c}2,3,4\end{subarray}}$ can be acquired via human parsing [60, 56], where $\mathcal{P}_{\begin{subarray}{c}2\end{subarray}}$ contains the head information, $\mathcal{P}_{\begin{subarray}{c}3\end{subarray}}$ masks the try-on related area using large areas of gray pixels, and $\mathcal{P}_{\begin{subarray}{c}4\end{subarray}}$ deletes the masked area with background color. Comparatively, $\mathcal{P}_{\begin{subarray}{c}3\end{subarray}}$ contains some clues about pose and shape while $\mathcal{P}_{\begin{subarray}{c}4\end{subarray}}$ completely removes the try-on related body area.
@@
Landmark $\mathcal{P}_{\begin{subarray}{c}13\end{subarray}}$ provides explicit semantic information for shape alignment. It guides clothing deformation by constraining the distance between clothing landmarks and corresponding body landmarks. SAL-VTON[77] makes improvements based on HR-Net[79] and proposes a Landmark estimation network. To train this network, a open source landmark dataset has been annotated on three mainstream datasets: VITON[21], VITON-HD[39], and DressCode[80]. The landmark shown in the Fig. 4 are derived from [49] but have not been made publicly available yet.
The choices of representative methods are shown in Table I, where pose keypoints are most commonly used and silhouette falls behind the other types. Additionally, there is no obvious preference towards representation types and a combination of several types could also be used.
@@ 2.4 Try-On Indication
Given the cloth-agnostic person representation and target clothing information as input, the Try-On Indication network (i.e., the orange module in Fig. 2) is used to predict the dressed person representation, which directly influences the generation of the final try-on image. The architectures of this module and corresponding constraint losses for representative methods are shown in Table I, where the encoder-decoder structure is the mainstream framework for Try-On Indication. It encodes cloth-agnostic person representation and decodes the representation of dressed person under the condition of the target clothing.
@@
As shown in Table I, the commonly-used loss for constraining Try-On Indication module mainly involve mask L1/L2, cross entropy loss, adversarial loss and perceptual loss. Specifically, mask L1/L2 is used to constrain the clothing mask, cross entropy loss facilitates the segmentation prediction and perceptual loss is designed for the generated RGB image. Adversarial loss is adopted for GAN-based generation methods. Both of Focal loss [65] and LSGAN loss [68] are designed to constrain the generation of segmentation.
@@ 2.5 Cloth Warping
Shown as the yellow module in the pipeline (Fig. 2, Cloth Warping module aims to transform the spatial distribution of the clothing image/feature to match the body. The mainstream deformation methods can be classified into: Thin Plate Spline (TPS), Spatial Transformation Network (STN), Flow Net and Implicit transformation. The choices of representative methods and corresponding losses are illustrated in Table I.
(a) TPS
(b) STN
(c) Flow
(d) Number Of Works / Year
Figure 5: Illustration for several spatial transformation approaches: (a) Thin Plate Spline (TPS), (b) Spatial Transformation Network (STN) and (c) Flow. (d) shows the usage statistics of various methods obtained after analyzing the literature listed in Table I and [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129].
@@ 2.5.1 Thin Plate Spline
Thin Plate Spline (TPS) interpolation method use the deformation of a thin steel plate to simulate 2D deformation, which is utilized for warping clothes in the literature[21, 23, 25, 22, 24, 29, 27, 28, 39, 73, 70]. Suppose there are N control points (e.g., the nodes of image grid shown in Fig. 5a), we denote the original coordinates as $T={(t_{1},t_{2},\ldots\ldots,t_{n})}^{T}$ and target coordinates as $Y={(y_{1},y_{2},\ldots\ldots,y_{n})}^{T}$. Take one point $y_{i}$ as an example, its interpolation under TPS transformation can be formulated as:
@@
Given a series of control points and their target positions, target image under TPS interpolation could be computed. However, for image-based virtual try-on, the lack of ground truth for target positions becomes an obstacle for TPS-based methods. VITON utilize shape context matching [130] to estimate target positions based on the context relationship of the clothing mask and the predicted warped mask. WUTON[36] uses TPS transformation at the feature level, which can enhance the diversity of clothing warping but has a shortage of image clarity. CP-VTON [23] designs a Geometry Matching Module (GMM, one type of Spatial Transformer Network) to feed $\mathcal{P}_{\begin{subarray}{c}2,5,6\end{subarray}}$ and learns TPS parameters for clothing warping. CP-VTON uses pixel-level L1 loss of warped clothing image to train GMM.
With the help of spatial transformation network (STN, which will be introduced next), further trials have been made to better constrain TPS transformation: 1) Regularization. To constrain the warping of clothing, CP-VTON+[22] introduces a regularization term that calculates the sum of distance changes for N nodes of grid; LA-VTON[29] proposes the constraint of grid interval consistency, which limits the distance between adjacent nodes; ACGPN[26] proposes a second-order difference constraint to prevent exaggerated deformation; DCTON[71] uses a homography matrix $H$ to reduce the change of the transformation matrix. 2) Auxiliary information. LM-VTON[24] designs a landmark-based method for fine-grained constraint. VITON-HD[39] additionally uses $\mathcal{P}_{\begin{subarray}{c}3,5\end{subarray}}$ and the clothing region Mask, while C-VTON[73] uses Densepose features alone as the input of GMM. 3) New model. CIT[70] replaces the association layer in GMM with Transformer structure to better calculate the spatial correlation between the person pose and the clothing.
@@ 2.5.2 Spatial Transformation Network
Spatial Transformation Network (STN) [62] can be embedded into a layer of the network, and consists of three main parts as shown in Fig. 5b: 1) Localisation Network is a simple regression network which contains several convolutional layers to regress $N$ variables (for example, a 2$\ast$3 matrix with 6 variables in the case of affine transformation). 2) Grid Generator is responsible for calculating the coordinates of the target image corresponding to the coordinates of the original image through matrix operations with $T_{\theta}(G)$. 3) Sampler samples the original image according to the coordinates in $T_{\theta}(G)$ and copies the pixels from original image to target image.
@@ 2.5.3 Flow Estimation
Dior[35] employs a single-layer flownet to deform the clothing features, where GFFE (Global Flow Field Estimator) proposed by GFLA[131] is used to predict flow map and soft mask. To capture more 3D changes of human posture deformation, 3D-GCL[75] uses [63] to create SMPL[132] flows as ground truth to constrain flow map, which results in better visual outcomes on the Deep Fashion data set.
@@ 2.5.4 Implicit Transformation
As shown in Fig. 5d, there has been a diversification in clothing warping methods in recent years. Single-STN methods are unable to handle natural deformation. Apart from a few works employing Multi-STN, most methods use a combination of STN and TPS, and they are classified under the TPS category. Influenced by the GMM framework, TPS transformation is the most popular method. TPS transformations have limitations in terms of degrees of freedom for deformation. In order to achieve more diverse clothing warping results, flow estimation has become a breakthrough in deformation performance. At the cost of greater computational overhead, flownet has achieved stronger deformation capabilities, and its usage is on the rise. Implicit transformation methods deform garments in a generative manner, further enhancing the diversity of garment deformation but also posing challenges in controlling clothing content. With the development of image generation technology under the diffusion framework, implicit transformation methods become promising.
Two kinds of loss for clothing warping are commonly adopted: 1) L1 loss and perceptual loss to supervise the warping with ground truth; 2) Regularization to alleviate exaggerated deformation, such as grid interval consistency, second-order-difference constraint, landmark loss, grid regularization, homography matrix regularization and TV (total variation) norm.
@@ 2.6 Try-On
Mask Combination: Some methods[21, 23, 25, 22, 70, 27, 28, 19] adopt the mask combination to generate the final try-on image, which is formulated as $I=I_{coarse}\cdot(1-M)+C\cdot M$. The input is the coarse try-on image $I_{coarse}$, the warped clothing image $C$ and the mask $M$ representing the clothing region of the dressed person. Essentially, the warped clothing is covered in the area represented by the $M$ in the $I_{coarse}$. This is a relatively simple and direct method, the advantage of which is that it will not increase the number of network parameters, but the disadvantage is that the quality of the final try-on image $I$ depends entirely on the correctness of the clothing warping and the coarse try-on image where misaligned regions cause artifacts.
Generation: Generation-based methods[26, 33, 24, 36, 34, 71, 37, 39, 35, 72, 48, 38, 49, 31, 64, 32, 29, 66, 74, 75] all use a generator to design the Try-On module. Among them, [36, 34, 72, 37, 38, 64, 30] use U-Net as the generator. Besides the try-on image as the generation target, VTNFP[30] uses U-Net to generate the try-on image and the combination mask at the same time, where combination mask is used to optimize the clothing details of the try-on image. The U-Net in ZFlow[72] additionally generates representation features $\mathcal{P}_{\begin{subarray}{c}8,9,10\end{subarray}}$ and calculates the loss between these three features and the ground truth to assist the training of the Try-On network.
@@
Previously, only using the Try-On module (without explicit try-on indication and cloth warping) could only generate rough try-on images. However, this issue has been alleviated in the era of large-scale models. TryOnDiffusion[50] has designed a substantial Parallel-UNet, which, after being trained with massive amounts of data, can directly use the original image and pose features to accomplish high-quality try-on.
As illustrated in Table I, besides the commonly-used L1 loss, perceptual loss and adversarial loss, some works further design local constraint (e.g., content preserving loss [71] and editing-localization loss [48]) and additional semantic constraint (e.g., feature matching loss [39] and attribute loss [49]).
3 Datasets and Evaluation Criteria
3.1 Datasets
In this section, we will introduce the currently public virtual try-on datasets. Before showing the details of each dataset, Table II compares them in terms of several key characteristics and counts the usage quantities of various datasets. VITON [21] is the first dataset for image-based virtual try-on and also the most popular one. Fig. 6 visually shows the samples for each dataset.
VITON dataset was released by Han et al.[21] in 2018, which collected a total of 16,253 pairs of top clothing images and frontal-view woman images wearing corresponding clothes. The clothing images were taken from the front with the clothing in a flat state and the background was pure white. The person body images were taken from the front upper body of the model, and the background was gray or light gray. 14,221 pairs of images are used as the train set, and 2,032 pairs of images are used as the test set.
TABLE II: Key properties of datasets. Popularity refers to the number of usage that is summarized from the literature listed in Table I and [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127].
@@ 3.1 Datasets
8
Figure 6: Examples in the dataset. VITON [21] contains pairs of female photo and corresponding in-shop clothing image. MPV[134] cosists of triplets of female photos in different poses with the same clothing and the corresponding clothing image. It also parses pose keypoints. Deep Fashion (In-Shop Clothes Retrieval Benchmark) [135] contains images of the same model in different poses. VITON-HD [39] is a high-resolution dataset accompanied with parsed segmentation, dense pose and pose keypoints. ESF [49] consists of clothed model photos for both male and female. Dress Code [80] involves more clothing types and provides dressed model photo, clothing image, pose keypoints, segmentation and dense pose.
MPV dataset was released by Dong et al.[134] in 2019, which contains 35,687 model images and 13,524 clothing images. The clothing images are the front images of the clothing in the flat state, and the model images are light gray background images. Each model in MPV corresponds to multiple images of different poses, and the image resolution is $256\times 192$. Dong et al. extracted 62,780 triplets of the same person wearing the same clothes but with different poses, further dividing 52,236 triplets as the training set, and shuffling 10,544 triplets as the test set.
@@
VITON-HD dataset was released by Choi et al.[39] in 2021, which comes from the same source of VITON, but its resolution is kept to $1024\times 768$. As shown in Fig. 6, it additionally contains parsing information such as segmentation, dense pose and pose keypoints. VITON-HD is divided into 11,647 image pairs for training and 2,032 image pairs for testing.
E-Shop Fashion(ESF) dataset was proposed by Feng et al.[49] in 2022, which contains 180,000 clothed model photos with adjusted resolution of $512\times 512$. Each image contains a front-standing pose model, and the model image is centrally aligned and filled to the region between the chin and thighs. Furthermore, the dataset is split into a training set containing 170,000 samples and a test set containing 10,000 samples.
Dress Code dataset was proposed by Morelli et al. in 2022 [80], which is more than three times larger than the publicly available virtual try-on datasets, with high-resolution paired images ($1024\times 768$). Different from VITON, which only contains upper-body female image pairs, Dress Code contains upper-body, lower-body, and full-body clothing for both female and male, as well as corresponding model try-on images. The images are divided into three categories: upper-body clothing (consisting of tops, T-shirts, shirts, sweatshirts and sweaters), lower-body clothing (consisting of skirts, pants, shorts, and leggings), and dresses. In total, the dataset consists of 53,795 image pairs: 15,366 pairs for upper-body clothing, 8,951 pairs for lower-body clothing, and 29,478 pairs for dresses. It is divided into a training set containing 48,392 pairs and a test set containing 5,400 pairs (1,800 pairs for upper body, lower body, and full body, respectively). The images are pre-processed and cut at the nose to ensure the privacy of the model. Dress Code also provides pose keypoints, dense pose, and segmentation mask for clothed model images.
In summary, although VITON is the most popular dataset, its limitations include a single type of clothing and a small amount of data. As a ground breaking work for image generation based virtual try-on, it could be used as the most basic benchmark with most works for comparison. However, it restricts the generalization performance. The most recently published Dress Code is a comprehensive dataset with the most diverse types of in-shop clothing images and the widest variety of models for trying on clothes, which should be a good choice for training virtual try-on models at present.

### v3

@@ 2.3 Cloth-agnostic Person Representation
RGB image $\mathcal{P}_{\begin{subarray}{c}1,2,3,4\end{subarray}}$ provides pixel-level features that can better preserve the identity information of individuals and the content outside the try-on area. $\mathcal{P}_{\begin{subarray}{c}1\end{subarray}}$ shows the original person image which is disturbed by the original clothes. $\mathcal{P}_{\begin{subarray}{c}2,3,4\end{subarray}}$ can be acquired via human parsing [13, 33, 132], where $\mathcal{P}_{\begin{subarray}{c}2\end{subarray}}$ contains the head information, $\mathcal{P}_{\begin{subarray}{c}3\end{subarray}}$ masks the try-on related area using large areas of gray pixels, and $\mathcal{P}_{\begin{subarray}{c}4\end{subarray}}$ deletes the masked area with background color. Comparatively, $\mathcal{P}_{\begin{subarray}{c}3\end{subarray}}$ contains some clues about pose and shape while $\mathcal{P}_{\begin{subarray}{c}4\end{subarray}}$ completely removes the try-on related body area.
@@
Landmark $\mathcal{P}_{\begin{subarray}{c}13\end{subarray}}$ provides explicit semantic information for shape alignment. It guides clothing deformation by constraining the distance between clothing landmarks and corresponding body landmarks. SAL-VTON [128] makes improvements based on HR-Net [114] and proposes a landmark estimation network.
The choices of representative methods are shown in Table 1, where pose keypoints are most commonly used and silhouette falls behind the other types. Additionally, there is no obvious preference towards representation types and a combination of several types could also be used.
@@ 2.4 Try-On Indication
Given the cloth-agnostic person representation and target clothing information as input, the Try-On Indication network (i.e., the orange module in Fig. 2) is used to predict the dressed person representation, which directly influences the generation of the final try-on image. The architectures of this module and corresponding constraint losses for representative methods are shown in Table 1, where the encoder-decoder structure is the mainstream framework for Try-On Indication. It encodes cloth-agnostic person representation and decodes the representation of dressed person under the condition of the target clothing.
@@
As shown in Table 1, the commonly-used loss for constraining Try-On Indication module mainly involve mask L1/L2, cross entropy loss, adversarial loss and perceptual loss. Specifically, mask L1/L2 is used to constrain the clothing mask, cross entropy loss facilitates the segmentation prediction and perceptual loss is designed for the generated RGB image. Adversarial loss is adopted for GAN-based generation methods. Both of Focal loss [74] and LSGAN loss [81] are designed to constrain the generation of segmentation.
@@ 2.5 Cloth Warping
Shown as the yellow module in the pipeline (Fig. 2, Cloth Warping module aims to transform the spatial distribution of the clothing image/feature to match the body. The mainstream deformation methods can be classified into: Thin Plate Spline (TPS), Spatial Transformation Network (STN), Flow Net and Implicit transformation. The choices of representative methods and corresponding losses are illustrated in Table 1.
Figure 5: Illustration for several spatial transformation approaches: (a) Thin Plate Spline (TPS), (b) Spatial Transformation Network (STN) and (c) Flow. (d) shows the usage statistics of existing methods.
@@ 2.5.1 Thin Plate Spline
Thin Plate Spline (TPS) interpolation method use the deformation of a thin steel plate to simulate 2D deformation, which is utilized by most methods for warping clothes. Suppose there are N control points (e.g., the nodes of image grid shown in Fig. 5(a)), we denote the original coordinates as $T={(t_{1},t_{2},\ldots\ldots,t_{n})}^{T}$ and target coordinates as $Y={(y_{1},y_{2},\ldots\ldots,y_{n})}^{T}$. Take one point $y_{i}$ as an example, its interpolation under TPS transformation can be formulated as:
@@
Given a series of control points and their target positions, target image under TPS interpolation could be computed. However, for image-based virtual try-on, the lack of ground truth for target positions becomes an obstacle for TPS-based methods. VITON [37] utilizes shape context matching [7] to estimate target positions based on the context relationship of the clothing mask and the predicted warped mask. CP-VTON [119] designs a Geometry Matching Module (GMM, one type of Spatial Transformer Network) to feed $\mathcal{P}_{\begin{subarray}{c}2,5,6\end{subarray}}$ and learns TPS parameters for clothing warping. WUTON [47] uses TPS transformation at the feature level, which can enhance the diversity of clothing warping but has a shortage of image clarity.
With the help of spatial transformation network (STN, which will be introduced next), further trials have been made to better constrain TPS transformation: 1) Regularization. To constrain the warping of clothing, LA-VTON [60] proposes the constraint of grid interval consistency, which limits the distance between adjacent nodes; CP-VTON+ [82] introduces a regularization term that calculates the sum of distance changes for N nodes of grid; ACGPN [130] proposes a second-order difference constraint to prevent exaggerated deformation; DCTON [30] uses a homography matrix $H$ to reduce the change of the transformation matrix. 2) Auxiliary information. LM-VTON [75] designs a landmark-based method for fine-grained constraint. VITON-HD [18] additionally uses $\mathcal{P}_{\begin{subarray}{c}3,5\end{subarray}}$ and the clothing region Mask, while C-VTON [26] uses Densepose features alone as the input of GMM. 3) New model. CIT [97] replaces the association layer in GMM with Transformer structure to better calculate the spatial correlation between the person pose and the clothing.
@@ 2.5.2 Spatial Transformation Network
Spatial Transformation Network (STN) [48] can be embedded into a layer of the network, and consists of three main parts as shown in Fig. 5(b): 1) Localisation Network is a simple regression network which contains several convolutional layers to regress $N$ variables (for example, a $2\times 3$ matrix with 6 variables in the case of affine transformation). 2) Grid Generator is responsible for calculating the coordinates of the target image corresponding to the coordinates of the original image through matrix operations with $T_{\theta}(G)$. 3) Sampler samples the original image according to the coordinates in $T_{\theta}(G)$ and copies the pixels from original image to target image.
@@ 2.5.3 Flow Estimation
Dior [22] employs a single-layer flownet to deform the clothing features, where GFFE (Global Flow Field Estimator) proposed by [99] is used to predict flow map and soft mask. To capture more 3D changes of human posture deformation, 3D-GCL [46] uses flownet [68] to create SMPL [78] flows as ground truth to constrain the flow map, which results in better visual outcomes on the Deep Fashion dataset.
@@ 2.5.4 Implicit Transformation
As shown in Fig. 5(d), there has been a diversification in clothing warping methods in recent years. Single-STN methods are unable to handle natural deformation. Apart from a few works employing Multi-STN, most methods use a combination of STN and TPS, and they are classified under the TPS category. Influenced by the GMM framework, TPS transformation is the most popular method. TPS transformations have limitations in terms of degrees of freedom for deformation. In order to achieve more diverse clothing warping results, flow estimation has become a breakthrough in deformation performance. At the cost of greater computational overhead, flownet has achieved stronger deformation capabilities, and its usage is on the rise. Implicit transformation methods deform garments in a generative manner, further enhancing the diversity of garment deformation but also posing challenges in controlling clothing content. With the development of image generation technology under the diffusion framework, implicit transformation methods become promising, whereas diffusion based methods usually directly synthesize the final try-on image without explicitly warping the clothes in advance.
Two kinds of loss for clothing warping are commonly adopted: 1) L1 loss and perceptual loss to supervise the warping with ground truth; 2) Regularization to alleviate exaggerated deformation, such as grid interval consistency [60], second-order-difference constraint [130, 18, 126], landmark loss [75, 128], grid regularization [82, 97], homography matrix regularization [30] and TV (total variation) norm [38, 20, 61].
@@ 2.6 Try-On
Mask Combination: Such methods adopt the mask combination to generate the final try-on image, which is formulated as $I=I_{coarse}\cdot(1-M)+C\cdot M$. The input is the coarse try-on image $I_{coarse}$, the warped clothing image $C$ and the mask $M$ representing the clothing region of the dressed person. Essentially, the warped clothing is covered in the area represented by the $M$ in the $I_{coarse}$. This is a relatively simple and direct method, the advantage of which is that it will not increase the number of network parameters, but the disadvantage is that the quality of the final try-on image $I$ depends entirely on the correctness of the clothing warping and the coarse try-on image where misaligned regions cause artifacts.
Generation: Generation-based methods all use a generator to design the Try-On module. Among them, [47, 65, 20, 31, 40, 96, 135] use U-Net as the generator. Besides the try-on image as the generation target, VTNFP [135] uses U-Net to generate the try-on image and the combination mask at the same time, where combination mask is used to optimize the clothing details of the try-on image. The U-Net in ZFlow [20] additionally generates representation features $\mathcal{P}_{\begin{subarray}{c}8,9,10\end{subarray}}$ and calculates the loss between these three features and the ground truth to assist the training of the Try-On network.
@@
Table 2: Key properties of datasets. Popularity refers to the number of usage that is summarized from existing methods.
@@
11
Previously, only using the Try-On module (without explicit try-on indication and cloth warping) could only generate rough try-on images. However, this issue has been alleviated in the era of large-scale models. TryOnDiffusion [142] has designed a substantial Parallel-UNet, which, after being trained with massive amounts of data, can directly use the original image and pose features to accomplish high-quality try-on. More recent diffusion based methods put additional emphasis on attention mechanism [56] and acceleration strategy [136].
As illustrated in Table 1, besides the commonly-used L1 loss, perceptual loss and adversarial loss, some works further design local constraint (e.g., content preserving loss [30] and editing-localization loss [62]) and additional semantic constraint (e.g., feature matching loss [18] and attribute loss [27]). Diffusion-based methods usually adopt denoising loss, and [56] additionally introduces attention total variation loss.
3 Datasets and Evaluation Criteria
3.1 Datasets
In this section, we will introduce the currently public virtual try-on datasets. Before showing the details of each dataset, Table 2 compares them in terms of several key characteristics and counts the usage quantities of various datasets. VITON [37] is the first dataset for image-based virtual try-on and also the most popular one. Fig. 6 visually shows the samples for each dataset.
VITON dataset was released by [37], which collected a total of 16,253 pairs of top clothing images and frontal-view woman images wearing corresponding clothes. The clothing images were taken from the front with the clothing in a flat state and the background was pure white. The person body images were taken from the front upper body of the model, and the background was gray or light gray. 14,221 pairs of images are used as the train set, and 2,032 pairs of images are used as the test set.
Figure 6: Examples in the dataset. VITON contains pairs of female photo and corresponding in-shop clothing image. MPV cosists of triplets of female photos in different poses with the same clothing and the corresponding clothing image. It also parses pose keypoints. Deep Fashion (In-Shop Clothes Retrieval Benchmark) contains images of the same model in different poses. VITON-HD is a high-resolution dataset accompanied with parsed segmentation, dense pose and pose keypoints. ESF consists of clothed model photos for both male and female. Dress Code involves more clothing types and provides dressed model photo, clothing image, pose keypoints, segmentation and dense pose.
MPV dataset was released by [23], which contains 35,687 model images and 13,524 clothing images. The clothing images are the front images of the clothing in the flat state, and the model images are light gray background images. Each model in MPV corresponds to multiple images of different poses, and the image resolution is $256\times 192$. Dong et al. extracted 62,780 triplets of the same person wearing the same clothes but with different poses, further dividing 52,236 triplets as the training set, and shuffling 10,544 triplets as the test set.
@@ 3.1 Datasets
VITON-HD dataset was released by [18], with a resolution of $1024\times 768$. As shown in Fig. 6, it additionally contains parsing information such as segmentation, dense pose and pose keypoints. VITON-HD is divided into 11,647 image pairs for training and 2,032 image pairs for testing.
E-Shop Fashion(ESF) dataset was proposed by [27], which contains 180,000 clothed model photos with adjusted resolution of $512\times 512$. Each image contains a front-standing pose model, and the model image is centrally aligned and filled to the region between the chin and thighs. Furthermore, the dataset is split into a training set containing 170,000 samples and a test set containing 10,000 samples.
Dress Code dataset was proposed by [83], which is more than three times larger than the publicly available virtual try-on datasets, with high-resolution paired images ($1024\times 768$). Different from VITON, which only contains upper-body female image pairs, Dress Code contains upper-body, lower-body, and full-body clothing for both female and male, as well as corresponding model try-on images. The images are divided into three categories: upper-body clothing (consisting of tops, T-shirts, shirts, sweatshirts and sweaters), lower-body clothing (consisting of skirts, pants, shorts, and leggings), and dresses. In total, the dataset consists of 53,795 image pairs: 15,366 pairs for upper-body clothing, 8,951 pairs for lower-body clothing, and 29,478 pairs for dresses. It is divided into a training set containing 48,392 pairs and a test set containing 5,400 pairs (1,800 pairs for upper body, lower body, and full body, respectively). The images are pre-processed and cut at the nose to ensure the privacy of the model. Dress Code also provides pose keypoints, dense pose, and segmentation mask for clothed model images.
In summary, although VITON is the most popular dataset, its limitations include a single type of clothing and a small amount of data. As a ground breaking work for image generation based virtual try-on, it could be used as the most basic benchmark with most works for comparison. However, it restricts the generalization performance. The most recently published Dress Code is a comprehensive dataset with the most diverse types of in-shop clothing images and the widest variety of models for trying on clothes, which could be a good choice for training virtual try-on models at present.